# Laser Interactions with Gas Jets: EMP Emission and Nozzle Damage

### v2

@@ Abstract
We present a theoretical model of electromagnetic pulse emission and nozzle damage from high-power laser interactions with gas jets. As hot electrons are accelerated beyond the laser channel, they produce a positively-charged plasma that expands into the ambient gas until it reaches the gas jet nozzle. A strong plasma potential leads to electrical breakdown that ionises the gas even faster than the expanding plasma ions. When contact is made between the plasma and the conducting nozzle surface, a discharge current propagates to ground and an electromagnetic pulse is radiated via an antenna emission mechanism. Damage to gas jet nozzles is caused by heating from plasma ions rather than Ohmic heating via a discharge current. Our model agrees with experimental data from experiments on proton acceleration in near-critical plasmas at the Vulcan and VEGA-3 lasers.
May 1, 2024
@@ I Introduction
A second important concern of the laser-gas interaction is damage to the gas jet nozzle. Many of the most exciting applications of laser-gas research rely on high shot rates and a reproducible gas density profile to generate bright, high-fluence sources of energetic ions [11, 12, 13, 14, 15, 16], electrons[17, 18] or x-rays [19, 20]. Repeated melting of the gas nozzle therefore represents a serious and expensive hindrance to this research. On energetic systems like the Vulcan Petawatt laser [21, 22] at the Rutherford Appleton Laboratory, a single shot on a dense gas target is sufficient to destroy the gas nozzle completely (see Fig. 1). For systems operating at lower energy, lower gas density and shorter pulse duration, nozzle damage is more progressive but still leads to significant smoothing of the gas density profile [23, 24, 25], degradation of the laser interaction and reduced data reproducibility. Understanding the nozzle damage mechanism is therefore important for future laser-gas applications.
@@
To date, no theoretical framework has been proposed for the emission of EMP from gas jets and this makes it difficult to estimate the severity of its impact on new laser systems. In Sec. II and Sec. III.1-III.3, we present a model of laser-plasma expansion in a gas, where ionisation is mainly caused by ions streaming from the laser focus (anode) to the nozzle (cathode). In Sec. III.4, however, we show that electrical breakdown induced by a strong plasma-nozzle potential can fully ionise the gas on shorter timescales of 10-20 ps. EMP is then emitted similarly to solid-target interactions, where a discharge current propagates to ground (the chamber) along a cm-scale antenna (the nozzle or target support), emitting radiation at gigahertz frequencies. According to our model, the EMP amplitude for Vulcan Petawatt and VEGA-style interactions can reach $\sim 100$ $\mu$T or $\sim 100$ kV/m at a distance $\sim 1$ m from the target. The theoretical model is supported in Sec. VI by the results from two experiments on the VEGA-3 laser, which show reasonable agreement with the discharge time, magnetic field spatial profile and accumulated target charge.
In Sec. V, we consider how the nozzle may be damaged using the plasma expansion model. Two mechanisms are considered: (i) a kA-level discharge current and (ii) ion collisional heating. We show that damage to gas jet nozzles is more likely caused by the impact of plasma ions than Ohmic heating of the nozzle surface. Finally, in Sec. VII, we present ideas for how the EMP emission model can be reliably benchmarked with simulations and dedicated experiments. The practical impact of this work is broad: allowing scientists to reduce damage to expensive gas jet nozzles and minimize the electrical disruption of equipment. It likewise represents a rich seam of more fundamental research, connecting the physics of laser-target charging, ionisation, ion acceleration, high-voltage breakdown and antenna emission.
@@ II Expansion of a Laser-Plasma in a Gas Jet
We consider an expanding plasma created as a laser pulse propagates through a high-density gas. The laser pulse ionises the gas and creates a plasma channel. Plasma electrons are heated by the laser radiation to relativistic energies, and some of them escape the channel and leave it positively charged. This charged plasma cylinder then expands into the surrounding gas or plasma, depending on the efficiency of the ionisation mechanisms discussed in Sec. III. If the plasma expands into gas, the gas is ionised until contact is made with the conducting tip of the gas nozzle. Once contact is made and the plasma is connected directly to ground, a discharge is triggered and EMP radiation is emitted. Nozzle damage is determined variously by the amount of energy stored in the plasma, the strength of the discharge and the nozzle material.
Starting from some characteristic gas and laser parameters, we first explore the physics of plasma expansion, electrical discharge and EMP emission. By way of example, we consider the Vulcan-Petawatt laser, with pulse energy $E_{L}=200$ J, duration $\tau_{L}=600$ fs and wavelength $\lambda_{L}=1.053\,\mu$m [26]. The laser beam is focused using $f_{\#}=3$ optics to a spot with full width at half maximum (FWHM) $w_{L}=3.9\,\mu$m. Only 40% of the total laser energy is contained within the FWHM, for a maximum on-target intensity of $I_{L}\simeq 2.9\times 10^{20}$ W cm-2. The spatial extent of the laser channel in the gas depends on the wavelength of the laser and the gas density. For typical gas pressures $p_{\rm gas}$ ranging from 0.03 to 30 bar, the atomic/molecular densities in the laser focal region are in the range $n_{\rm gas}\sim 10^{18}-10^{21}$ cm-3. The maximum electron density in a fully ionised plasma is smaller than the critical density, so the laser can propagate through the gas.
The Rayleigh length of a diffraction-limited Gaussian beam in our example is $Z_{R}=\pi w_{L}^{2}/\lambda_{L}\simeq 44\,\mu$m. Since the gas jet width is usually larger than $Z_{R}$ and the beam is susceptible to relativistic self-guiding, the laser produces a channel a few times longer than $Z_{R}$ that can be viewed as a $l_{p}\simeq 0.5$ mm-long plasma cylinder created at a typical height of $h\simeq 0.5$ mm above the nozzle. We take the initial radius of the plasma channel to be a factor of a few times larger than the laser focal radius, $r_{p}=20\,\mu$m, to reflect the average size of the channel as the laser deposits energy over the full length of its path through the gas. We assume that laser absorption in the gas jet is $\eta_{L}\sim 90$%, with 40% transferred to hot electrons. This is consistent both with Particle-in-Cell (PIC) simulations[26, 27] and experimental interferograms [28], which show that a high-power laser pulse deposits a large fraction of its energy in the gas jet and leaves behind it a plasma channel several tens of microns wide.
Parameter Vulcan-TAP VEGA-3
Gas H2 He
$\lambda_{L}$ [$\mu$m] 1.05 0.8
$E_{L}$ [J] 80 8
$\tau_{L}$ [fs] 600 75
$f_{\#}$ 3 10
$w_{L}$ [$\mu$m] 3.9 9.8
$I_{L}$ [W cm-2] $2.9\times 10^{20}$ $3.6\times 10^{19}$
$a_{0}$ 15 4.1
$\eta_{L}$ 0.9 0.9
$\eta_{\rm las\to h}$ 0.4 0.1
$p_{\rm gas}$ [bar] 15 4.5
$h$ [$\mu$m] 500 500
$r_{p}$ [$\mu$m] 20 20
$V_{p}$ [cm3] $6\times 10^{-7}$ $6\times 10^{-7}$
$n_{e}$ [cm-3] $8.1\times 10^{20}$ $2.4\times 10^{20}$
$T_{e}$ [keV] 280 140
$T_{h}$ [MeV] 5.0 1.1
$\Phi_{p}$ [MV] 25 4.4
$Q$ [nC] 230 42
Table 1: Representative experimental parameters - as described in the text and Refs. [26, 24] - and estimates of plasma characteristics according to the model presented in Sec. II. The plasma radius $r_{p}$ and the laser energy fraction converted to hot electrons $\eta_{\rm las\to h}$ are estimated from dedicated PIC simulations.
The parameters of the experiments considered in this paper are given in Tab. 1 along with predictions from our model. Depending on the nozzle design and envisioned application, the laser pulse can be sent parallel to the surface of the nozzle tip at a height varying from a few tenths of a millimetre to a few millimetres. The width of the laser channel also depends on the laser focusing optics, laser power, and gas density profile along the laser axis. Figure 2 gives a schematic overview of the situation.
@@
II.1 Electric Charging of the Plasma
@@
where $C_{e}(T_{e})=3-[1-K_{1}(1/\Theta_{e})/K_{2}(1/\Theta_{e})]/\Theta_{e}$ is the heat capacitance of a relativistic electron gas. For the parameters of the Vulcan experiment given in Tab. 1, a plasma channel of radius $r_{p}=20\,\mu$m and length $l_{p}=0.5$ mm has volume $V_{p}=\pi r_{p}^{2}l_{p}\simeq 6\times 10^{-7}$ cm3. A hydrogen gas pressure of 15 bar corresponds to an atomic gas density $n_{\rm at}\simeq 8.1\times 10^{20}$ cm-3 at a distance $h\simeq 0.5$ mm above the nozzle, so we have $N_{e}\simeq 5.1\times 10^{14}$ electrons. Equation (2) yields an effective electron temperature $T_{e}=250$ keV for the total bulk energy of 43 J.
To estimate the hot electron density, we prescribe a hot electron temperature of $T_{h}=5.0$ MeV following the ponderomotive scaling[34], $(\sqrt{1+a_{0}^{2}}-1)m_{e}c^{2}$. The hot electron density can then be defined knowing the fraction of energy transferred to hot electrons $\eta_{\rm las\to h}$:
@@
Knowing the hot electron temperature and assuming that $\eta_{\rm las\to h}=0.4$ on Vulcan, this equation gives a hot electron fraction $n_{h}/n_{e}\simeq 2.5$% with total energy 29 J.
@@
where $\omega_{ph}=(e^{2}n_{h}/(m_{e}\epsilon_{0}))^{1/2}$ is the hot electron plasma frequency. The solution of this equation for typical parameters $h/r_{p}=250$ and $r_{p}\omega_{ph}/c=18$ gives $\Phi_{p}=24.8$ MV. Knowing the plasma potential and the plasma capacitance $C\simeq 0.0086$ pF, we find the charge ejected from the plasma column, $Q=C\Phi_{p}\simeq 234$ nC. This corresponds to a fraction of escaped electrons $\delta n_{h}/n_{e}\approx 2.8\times 10^{-3}$ and is consistent with the experiment [36]. Since the time required for a hot electron to cross the channel is relatively short ($\sim 2r_{p}/c=0.13$ ps), the time of plasma charging is determined by the propagation time of the laser pulse through the gas, $\sim 3$ ps.
The total energy of escaped electrons
@@
is split between the electrostatic energy of the charged plasma $\mathcal{E}_{\rm es}=Q^{2}/2C\simeq 3.2$ J and the kinetic energy of the escaped electrons $\simeq 4.2$ J.
II.2 Ion Acceleration in Expanding Plasma
Following a short phase of gas ionisation, electron heating and plasma charging, the plasma expands and cools down. The time required for the plasma to expand from the laser-generated channel to the nozzle defines the plasma discharge time. In Sec. IV, we show that the discharge time determines the energy and amplitude of the radiated EMP. Plasma x-ray emission and laser-accelerated fast electrons and ions, as shown in Sec. III, partially ionise the gas and provide seed electrons for a discharge via electrical breakdown.
@@ II.2 Ion Acceleration in Expanding Plasma
After the end of the laser pulse, there is no more energy supply and the plasma expands adiabatically under the electron thermal pressure. For the plasma parameters given above, the pressure $p_{e}=n_{e}T_{e}\simeq 330$ Mbar is much higher than the ambient gas pressure, so the plasma expands freely with electrons transferring their energy to ions. The plasma charge is conserved during the expansion phase and the potential decreases logarithmically as the plasma radius increases (see Eq. (5)). The plasma expansion can be described in some special cases by a self-similar rarefaction wave model [37, 38]. More detailed analysis is performed for a spherical plasma expansion in Ref. [39]. The ion energy distribution depends on the density profile and on the ratio of the plasma radius to the Debye length. For our purposes, it is sufficient to estimate the average ion kinetic energy by dividing the deposited laser energy by the number of ions in the plasma. For a hydrogen plasma, it is 1.5 times the bulk electron temperature $\varepsilon_{i}\simeq 0.4$ MeV. This corresponds to an ion velocity of $v_{i}\simeq 9\,\mu$m/ps, and an ion expansion time to the nozzle $t_{\rm exp}\simeq 60$ ps. The total energy carried by these ions is of order the bulk electron energy - that is, about $(1-\eta_{\rm las\to h})\eta_{L}E_{L}\sim 40$ J for the Vulcan experiment.

### v3

@@ Abstract
Understanding the physics of electromagnetic pulse emission and nozzle damage is critical for the long-term operation of laser experiments with gas targets, particularly at facilities looking to produce stable sources of radiation at high repetition rate. We present a theoretical model of plasma formation and electrostatic charging when high-power lasers are focused inside gases. The model can be used to estimate the amplitude of gigahertz electromagnetic pulses (EMPs) produced by the laser and the extent of damage to the gas jet nozzle. Looking at a range of laser and target properties relevant to existing high-power laser systems, we find that EMP fields of tens to hundreds of kV/m can be generated several metres from the gas jet. Model predictions are compared with measurements of EMP, plasma formation and nozzle damage from two experiments on the VEGA-3 laser and one experiment on the Vulcan Petawatt laser.
October 9, 2024
@@ I Introduction
A second important concern of the laser-gas interaction is damage to the gas jet nozzle. Many of the most exciting applications of laser-gas research rely on high shot rates and a reproducible gas density profile to generate bright, high-fluence sources of energetic ions [11, 12, 13, 14, 15, 16], electrons[17, 18] or x-rays [19, 20]. Repeated melting of the gas nozzle, therefore, represents a serious and expensive hindrance to this research. On energetic systems like the Vulcan Petawatt laser [21, 22] at the Rutherford Appleton Laboratory, a single shot on a dense gas target is sufficient to destroy the gas nozzle completely (see Fig. 1). For systems operating at lower energy, lower gas density and shorter pulse duration, nozzle damage is more progressive but still leads to significant smoothing of the gas density profile [23, 24, 25], degradation of the laser interaction and reduced data reproducibility. Understanding the nozzle damage mechanism is therefore important for future laser-gas applications.
@@
To date, no theoretical framework has been proposed for the emission of EMP from gas jets and this makes it difficult to estimate the severity of its impact on new laser systems. In Secs. II and III, we present a model of laser-plasma expansion in a gas, where ionisation is mainly caused by ions streaming from the laser focus (anode) to the nozzle (cathode). In Sec. III.4, however, we show that electrical breakdown induced by a strong plasma-nozzle potential can fully ionise the gas on shorter timescales of $10-20$ ps. EMP is then emitted similarly to solid-target interactions, where a discharge current propagates to ground (the chamber) along a cm-scale antenna (the nozzle or target support), emitting radiation at gigahertz frequencies. According to our model, the EMP amplitude for Vulcan Petawatt (Vulcan-TAP) and VEGA-3 interactions can reach tens to hundreds of $\mu$T or several tens of kV/m at a distance $\sim 1$ m from the target. The theoretical model is supported in Sec. VI by the results from two experiments on the VEGA-3 laser, which show reasonable agreement with the discharge time, magnetic field spatial profile and accumulated target charge.
In Sec. V, we consider how the nozzle may be damaged using the plasma expansion model. Two mechanisms are considered: (i) a kA-level discharge current and (ii) ion collisional heating. We show that damage to gas jet nozzles is more likely caused by the impact of plasma ions than by Ohmic heating of the nozzle surface. Model predictions for the EMP field strength and nozzle damage on different facilities is discussed in Sec. VII. Finally, in Sec. VIII, we present ideas for how the EMP emission model can be reliably benchmarked with simulations and dedicated experiments. The practical impact of this work is broad: allowing scientists to reduce damage to expensive gas jet nozzles and minimize the electrical disruption of equipment. It likewise represents a rich seam of more fundamental research, connecting the physics of laser-target charging, ionisation, ion acceleration, high-voltage breakdown and antenna emission.
@@ II Expansion of a Laser-Plasma in a Gas Jet
We consider an expanding plasma created as a laser pulse propagates through a high-density gas. The laser pulse ionises the gas and creates a plasma channel. Plasma electrons are heated by laser radiation to relativistic energies, and some of them escape the channel and leave it positively charged. This charged plasma cylinder then expands into the surrounding gas or plasma, depending on the efficiency of the ionisation mechanisms discussed in Sec. III. If the plasma expands into a gas, the gas is ionised until contact is made with the conducting tip of the gas nozzle. Once contact is made, and the plasma is connected directly to ground, a discharge is triggered and EMP radiation is emitted. Nozzle damage is determined variously by the amount of energy stored in the plasma, the strength of the discharge and the nozzle material.
Depending on the nozzle design and envisioned application, the laser pulse can be sent parallel to the surface of the nozzle tip at a height varying from a few tenths of a millimetre to a few millimetres. The width of the laser channel also depends on the laser focusing optics, laser power, and gas density profile along the laser axis. Figure 2 gives a schematic overview of the situation.
@@
II.1 Model of Electric Charging of the Plasma
The model of EMP emission proposed in this paper consists of a phase of plasma charging, followed by a discharge and antenna emission process. First, the plasma is formed, and the charge is lost as hot electrons escape the potential barrier. This charge supplies the electrostatic energy that is later radiated as EMP. The mechanism of EMP emission, where the plasma is discharged through the nozzle and radiates as a dipole antenna, is similar to the solid-target model already examined in various publications [3, 27, 2, 7].
@@ II.1 Model of Electric Charging of the Plasma
where $C_{e}(T_{e})=3-[1-K_{1}(1/\Theta_{e})/K_{2}(1/\Theta_{e})]/\Theta_{e}$ is the heat capacitance of a relativistic electron gas.
The hot electron density can then be defined knowing the fraction of energy transferred to hot electrons $\eta_{\rm las\to h}$:
@@
where $\omega_{ph}=(e^{2}n_{h}/m_{e}\epsilon_{0})^{1/2}$ is the hot electron plasma frequency.
The total energy of escaped electrons is split between the electrostatic energy of the charged plasma $\mathcal{E}_{\rm es}=Q^{2}/2C$ and the kinetic energy of the escaped electrons:
@@
II.2 Ion Acceleration in the Expanding Plasma
The number of ions in the laser plasma and their energy are important factors for determining the extent of damage to the gas nozzle, as well as for placing a lower limit on the speed of the plasma discharge. When the expanding thermal ions reach the nozzle surface they deposit their kinetic energy and can cause the nozzle to melt. Contact between the plasma and the nozzle also establishes an electrical path to ground, triggering a plasma discharge and the emission of EMP radiation.
The ion spectrum can be separated into two broad populations [34, 25]: (i) thermal ions that expand as part of the laser plasma and (ii) a less numerous population of fast ions accelerated in the charge separation field produced by the escaped hot electrons [35, 28]. We restrict ourselves to modelling the thermal ions, since they carry most of the energy.
@@ II.2 Ion Acceleration in the Expanding Plasma
Following a short phase of gas ionisation, electron heating and plasma charging, the plasma expands and cools down. After the end of the laser pulse, there is no more energy supply and the plasma expands adiabatically under the electron thermal pressure. The plasma pressure $p_{e}=n_{e}T_{e}$ is much higher than the ambient gas pressure, so the plasma expands freely with electrons transferring their energy to ions. The plasma charge is conserved during the expansion phase and the potential decreases logarithmically as the plasma radius increases (see Eq. (5)). The plasma expansion can be described in some special cases by a self-similar rarefaction wave model [36, 37]. More detailed analysis is performed for a spherical plasma expansion in Ref. [38]. In practice, the ion energy distribution depends on the density profile and on the ratio of the plasma radius to the Debye length. For our purposes, it is sufficient to estimate the average ion kinetic energy $\varepsilon_{i}$ by equating the ion and electron energy densities, $n_{i}\varepsilon_{i}\simeq n_{e}\varepsilon_{e}\approx\frac{3}{2}n_{e}T_{e}$, where $n_{i}$ is the ion density in the plasma channel and $\varepsilon_{e}$ is the average electron kinetic energy. The total number of ions in the plasma is $N_{i}=n_{i}V_{p}$.
II.3 Model estimates for Vulcan-TAP conditions
Laser Parameters
[…]
$I_{L}$ [W cm-2] $2.9\times 10^{20}$ $3.6\times 10^{19}$
$E_{L}$ [J] 80 8
$\tau_{L}$ [fs] 600 75
$w_{L}$ [$\mu$m] 3.9 9.8
$\lambda_{L}$ [$\mu$m] 1.05 0.8
$f_{\#}$ 3 10
$a_{0}$ 15 4.1
Table 1: Representative input laser parameters - measured or inferred - as described in the text and Refs. [26, 25].
Gas and Plasma Parameters (Measured)
[…]
Gas H2 He
$n_{e}$ [cm-3] $8.1\times 10^{20}$ $2.4\times 10^{20}$
$p_{\rm gas}$ [bar] 15 4.5
$h$ [$\mu$m] 500 500
Gas and Plasma Parameters (Assumed)
$r_{p}$ [$\mu$m] 20 20
$l_{p}$ [$\mu$m] 500 500
$\eta_{L}$ 1.0 0.9
$\eta_{\rm las\to h}$ 0.4 0.1
Table 2: Representative gas and plasma parameters (measured and assumed). The plasma radius $r_{p}$ and the laser energy fraction converted to hot electrons $\eta_{\rm las\to h}$ are estimated from dedicated PIC simulations[26, 25].
Plasma Properties
[…]
$V_{p}$ [cm3] $6\times 10^{-7}$ $6\times 10^{-7}$
$T_{e}$ [keV] 280 140
$T_{h}$ [MeV] 5.0 1.1
$\Phi_{p}$ [MV] 25 4.4
$Q$ [nC] 230 42
Table 3: Plasma properties estimated using the model presented in Sec. II.1.
The plasma charging model described in the previous section takes several laser and gas parameters as inputs which must either be experimentally determined or estimated by other means. Input parameters relevant to the experiments considered in this paper are given in Tabs. 1 and 2, along with the outputs from our model given in Tab. 3. We use Vulcan-TAP parameters in Secs. III and V to explore likely mechanisms of gas ionization and nozzle damage. The VEGA-3 parameters are used to benchmark our model against the experiments described in Sec. VI.
We assume the laser energy is separated as follows: the total laser energy is multiplied by some fraction to reflect the amount of energy contained within the laser focus, which is called $E_{L}$; the energy in the focus is then multiplied by a fraction $\eta_{L}$, which reflects its absorption in the gas and a further fraction $\eta_{\rm las\to h}$, which is the fraction of energy converted to hot electrons. The laser energy not converted into hot electrons goes into the thermal electrons and is eventually converted into the kinetic energy of the plasma ions as the plasma expands.
Four of the input parameters for our model are “assumed”, which means they have been estimated based on simulations or previous experimental data. The laser absorption $\eta_{L}$ in the gas is inferred from dedicated PIC simulations [25, 26], which indicate that the laser is entirely depleted after it has passed through the dense gas. The conversion efficiency of laser energy to hot electrons, $\eta_{\rm las\to h}$, is a function of many variables, including the gas density and laser intensity. It is not measured in any of the experiments described here and must therefore be estimated to the nearest order of magnitude. For near-critical density plasmas the available data is relatively scarce, though there is evidence that the hot electron conversion efficiency ranges from a few percent to more than ten percent under certain conditions [39, 40]. Here, we take an upper estimate of $\eta_{\rm las\to h}=0.4$ to illustrate a scenario of strong charging and EMP. The plasma channel radius and length are estimated from PIC simulations [25, 26, 31, 41] and experimental interferograms [25, 42], which suggest that channels several hundred microns in length and tens of microns in radius can be formed on the Vulcan-TAP and VEGA-3 laser systems a few ps after the arrival of the laser pulse.
Consider conditions on the Vulcan-Petawatt laser, with pulse energy $E_{\rm L,tot}=200$ J, duration $\tau_{L}=600$ fs and wavelength $\lambda_{L}=1.053\,\mu$m [26]. The laser beam is focused using $f_{\#}=3$ optics to a spot with full width at half maximum (FWHM) $w_{L}=3.9\,\mu$m. Only 40% of the total laser energy is contained within the FWHM, for a maximum on-target intensity of $I_{L}\simeq 2.9\times 10^{20}$ W cm-2. The spatial extent of the laser channel in the gas depends on the wavelength of the laser and the gas density. For typical gas pressures $p_{\rm gas}$ ranging from 0.03 to 30 bar, the atomic/molecular densities in the laser focal region are in the range $n_{\rm gas}\sim 10^{18}-10^{21}$ cm-3. The maximum electron density in a fully ionised plasma is smaller than the critical density, so the laser can propagate through the gas.
The Rayleigh length of a diffraction-limited Gaussian beam in our example is $Z_{R}=\pi w_{L}^{2}/\lambda_{L}\simeq 44\,\mu$m. Since the gas jet width is usually larger than $Z_{R}$ and the beam is susceptible to relativistic self-guiding, the laser produces a channel significantly longer [42, 43] than $Z_{R}$. It can be viewed as a $l_{p}\simeq 0.5$ mm-long plasma cylinder created at a typical height of $h\simeq 0.5$ mm above the nozzle. We take the initial radius of the plasma channel to be a factor of a few times larger than the laser focal radius, $r_{p}=20\,\mu$m, to reflect the average size of the channel as the laser deposits energy over the full length of its path through the gas. We assume that laser absorption in the gas jet is $\eta_{L}\sim 100$%, with 40% transferred to hot electrons.
For the parameters of the Vulcan experiment given in Tabs. 1 - 3, a plasma channel of radius $r_{p}=20\,\mu$m and length $l_{p}=0.5$ mm has volume $V_{p}=\pi r_{p}^{2}l_{p}\simeq 6\times 10^{-7}$ cm3. A hydrogen gas pressure of 15 bar corresponds to an atomic gas density $n_{\rm at}\simeq 8.1\times 10^{20}$ cm-3 at a distance $h\simeq 0.5$ mm above the nozzle, so we have $N_{e}\simeq 5\times 10^{14}$ electrons. Equation (2) yields an effective electron temperature $T_{e}=250$ keV for the total bulk energy of 43 J.
A hot electron temperature of $T_{h}=5.0$ MeV is prescribed following the ponderomotive scaling [44], $(\sqrt{1+a_{0}^{2}}-1)m_{e}c^{2}$. If we further assume that $\eta_{\rm las\to h}=0.4$ on Vulcan, then Eq. 4 gives a hot electron fraction $n_{h}/n_{e}\simeq 2.5$% with total energy 29 J. The hot electron density follows from the hot electron temperature and the heat capacity of a relativistic electron gas, as shown in the previous section.
The solution of Eq. (5) for typical parameters $h/r_{p}=25$ and $r_{p}\omega_{ph}/c=18$ gives $\Phi_{p}=24.8$ MV. Knowing the plasma potential and the plasma capacitance $C\simeq 0.0086$ pF, we find the charge ejected from the plasma column, $Q=C\Phi_{p}\simeq 234$ nC. This corresponds to a fraction of escaped electrons $\delta n_{h}/n_{e}\approx 2.8\times 10^{-3}$ and is consistent with the experiment [45].
Since the time required for a hot electron to cross the channel is relatively short $\sim 2r_{p}/c=0.13$ ps, the time of plasma charging is determined by the propagation time of the laser pulse through the gas, $\sim 3$ ps.
Following Eq. (6), the total energy of escaped electrons $\mathcal{E}_{\rm esc}$ is divided between the electrostatic energy of the charged plasma $\mathcal{E}_{\rm es}=Q^{2}/2C\simeq 3.2$ J and the kinetic energy of the escaped electrons $\simeq 4.2$ J.
As discussed in the previous section, the average ion kinetic energy, $\varepsilon_{i}\simeq 0.4$ MeV, is 1.5 times the bulk electron temperature for a hydrogen plasma. This corresponds to an ion velocity of $v_{i}\simeq 9\,\mu$m/ps, and an ion expansion time to the nozzle $t_{\rm exp}\simeq 60$ ps. The total energy carried by these ions is of order the energy carried by the expanding thermal electrons - that is, about $(1-\eta_{\rm las\to h})\eta_{L}E_{L}\sim 40$ J for the Vulcan experiment.
The model estimate of $5\times 10^{14}$ ions with a combined kinetic energy of several tens of joules can be compared with measured ion spectra and simulations. Fig. 3a shows an ion spectrum from Ref. [26] measured at $90^{\circ}$ to the laser axis. The peak ion yield of $\sim 10^{11}$ MeV-1sr-1 occurs at $\sim 2$ MeV. The proton signal drops off at lower energies because the configuration of B-/E-fields inside the spectrometer limits its dynamic range to about two decades. Other ion spectral measurements [46, 47, 34, 25, 16] and simulations [41, 26] suggest that the sub-MeV thermal ion population should be orders of magnitude more numerous than the multi-MeV plasma-accelerated ions. Ion spectra measured at the PHELIX laser using spectrometers specifically designed to measure sub-MeV ions [46] show sub-MeV ion yields of $\sim 10^{14}$ particles.
Figure 3: Ion spectra collected in gas jet experiments: (a) Proton spectrum from an experiment on Vulcan-TAP, measured at $90^{\circ}$ to the laser axis using a Thomson parabola spectrometer with BAS-TR image plate [26]. The shaded region is the $3\sigma$ detection limit. (b) $\alpha$-particle spectra from an experiment at VEGA-3 [25], measured at ±$\pm 17^{\circ}$ from the laser axis using Diamond time-of-flight detectors. The two spectra were recorded on different shots. See also Fig. 5a. (c) Proton spectrum from a separate experiment at VEGA-3 [23], measured using a Thomson parabola spectrometer at $90^{\circ}$ to the laser axis. The spectrometer dynamic range limits reliable measurements to energies $\gtrsim 1.3$ MeV. The blue dashed line represents the background noise level. See also Fig. 5b.